# PromptGAR: Flexible Promptive Group Activity Recognition

### v1

@@ 4.1 Experimental Setup
Volleyball Dataset [21] contains 3,493 clips for training and 1,337 clips for testing. Each clip has 41 frames and is labeled with one of eight group activities. Bounding box and skeleton annotations, provided separately by [21, 55], are available only for the central 16 frames of each clip. In approximately 20.5% of clips, bounding boxes are shifted substantially from their correct positions. Therefore we use skeleton-derived envelope bounding boxes instead. Skeleton annotations are also occasionally unreliable due to heavy occlusions or truncations in frame boundaries. Consistent with prior works, we limit our analysis for central 16 frames to avoid potential interference from additional group activities present in the remaining half of the clips.
Modalities Top1 Acc Mean Acc Method RGB Bbox Kpt Flow Ball InstID ARG [44] 90.7 91.0 HiGCIN [46] 91.4 92.0 DIN [50] $92.7^{\star}$ $92.8^{\star}$ POGARS [43] 93.9 - Composer [55] $93.6^{\star}$ SkeleTR [13] 94.4 - MP-GCN [32] $95.0^{\dagger}$ $95.0^{\dagger}$ CRM [4] 93.0 - ActorFormer [15] 94.4 - GIRN [37] 94.0 - SACRF [38] 95.0 - GroupFormer [29] 95.7 - Dual-AI [19] 95.4 - KRGFormer [36] 94.6 94.8 PromptGAR 96.0 96.3
Table 1: Quantitative Comparisons in Volleyball Dataset. (--) denotes not reported in the paper, (*) reproduced from released codes, ($\dagger$) reproduced with keypoint-ball results (late-fusion parts un-released), and unmarked other methods are not open-sourced.
Modalities Top1 Acc Mean Acc Method RGB Bbox Kpt Flow Ball InstID ARG [44] 59.0 56.8 ActorFormer [15] 47.1 41.5 SAM [47] 49.1 47.5 DIN [50] 61.6 56.0 Dual-AI [19] 58.1 50.2 KRGFormer [36] 72.4 67.1 DFWSGAR [24] 75.8 71.2 Flaming-Net [35] 79.1 76.0 MP-GCN [32] $75.8^{\dagger}$ $72.0^{\dagger}$ PromptGAR 80.6 76.9
@@
Modalities 2-Mosaic 4-Mosaic Method RGB Bbox Kpt InstID Area Re-train Top1 Mean Top1 Mean 93.7 94.1 89.9 89.3 89.6 90.2 81.5 82.7 PromptGAR 54.0 53.6 23.5 23.0
Table 3: Multi-Group Activity Recognition. Area prompts provide reliable results, empowering PromptGAR to selectively recognize group activities in multi-group settings. Experiments are conducted under smaller models with $O=16$.
NBA Dataset [47] includes 7,624 training clips and 1,548 testing clips. Each clip includes 72 frames and is categorized into nine group activity labels. Due to its increasing complexity, the NBA dataset demands special design and processing compared to the Volleyball dataset [21]. Unlike Volleyball [21], we use all 72 frames in NBA because key event frames are not centrally located. Also, NBA activities are much longer than Volleyball’s. For example, ’3p-fail-offensive’ requires recognizing shooter location, ball trajectory, and possession.
Mosaic Volleyball Dataset. We generated a multi-group activity recognition dataset from the Volleyball dataset, as existing datasets [21, 47] are limited to single-group scenarios. As depicted in Fig. 5, we created two types of multi-group videos: 2-Mosaic (two clips stacked vertically) and 4-Mosaic (four clips arranged in a $2\times 2$ grid). For each mosaic, the target clip is randomly assigned a position, and the remaining clips are uniformly selected from the Volleyball dataset. Annotations are derived from the target clip: its activity label is used for the entire mosaic, and its bounding boxes and skeleton data are adjusted based on its position. Area prompts are generated by creating a bounding envelope around all bounding boxes within each frame. We evaluate performance using Top-1 and Mean Accuracy.
@@ 4.2 Group Activity Recognition
Volleyball Dataset. In Tab. 1, we compare our method to the group activity recognition methods that rely on: (a) RGB frames, like DIN [50], which suffers from limited performance; (b) skeletons and ball trajectories, like Composer [55] and MP-GCN [32], which requires ball trajectories as extra inputs; (c) multi-modal techniques, like GroupFormer [29] and Dual-AI [19], which takes computationally expensive optical flows as inputs and thus sensitive to input frame rates. In contrast, our PromptGAR, without the drawbacks of those other modality inputs, still produces 96.0% top1 accuracy and 96.3% mean accuracy, achieving significant improvements across different baseline models.
NBA Dataset. Tab. 2 shows that PromptGAR significantly outperforms previous methods on the challenging NBA dataset [47]. This success is attributed to the instance ID modality, which provides stable long-term actor consistency, crucial for recognizing complex, multi-event group activities in longer videos.
Modalities Method RGB Bbox Kpt Flow InstID Re-train Top1 Acc Mean Acc 95.7 - 94.9 - GroupFormer [29] 94.1 - 96.0 96.3 95.7 95.8 94.1 94.3 PromptGAR 93.9 94.4
@@
4.3 Multi-Group Activity Recognition
As demonstrated in Tab. 3, even when provided with only the most basic ‘area prompts’ as minimal ‘selectors’ without any additional annotations, PromptGAR still achieved impressive recognition accuracy on multi-group mosaic videos. Specifically, it reaches top1 accuracy of 89.6% on the 2-Mosaic Volleyball dataset and 81.5% on the more complex 4-Mosaic Volleyball dataset. This demonstrates PromptGAR’s capability to effectively discern and recognize the particular group activity within multiple groups given the simple ‘area prompts’.
4.4 Input Flexibility
@@
4.5 Ablation Studies
Tab. 7 details the impact of our novel modules on performance, starting with a baseline of 96.0% top-1 accuracy on the Volleyball dataset. (a) Relative instance attention. Eliminating instance IDs during training and testing resulted in a 0.6% drop in top-1 accuracy. This confirms that instance IDs are vital for encoding long-term consistency, enhancing performance. Furthermore, relative instance attention also accelerated training convergence. (b) Prompt tokens in head. Removing prompt tokens $\widehat{\textbf{F}}_{p}$ from the head, leaving only the GAR class token $\widehat{\textbf{X}}_{gar}$, led to a 0.3% accuracy decrease. This highlights the necessity of explicitly incorporating prompt tokens. Without them, the model struggles to integrate multi-modal prompt features, as the GAR class token alone sufficiently captures RGB information. (c) No prompts. Disabling the entire prompt encoder, using only the GAR class token $\widehat{\textbf{X}}_{gar}$ as input for the recognition decoder and the head, further reduced accuracy by 0.8%. This underscores the contribution of multi-modal prompt inputs in PromptGAR. Crucially, even without any prompts during training and testing, our model still compares favorably with other prompt-based baseline methods. Specifically, it achieves a 1.6% higher accuracy than DIN, which utilizes both RGB frames and bounding boxes, and a 0.7% higher accuracy than Composer, which uses skeletons and ball trajectories as additional information. This demonstrates the superior ability of our MViTv2 architecture to process spatial-temporal information within only the RGB features compared to other backbones that process the video frame-by-frame separately.
@@ 4.5 Ablation Studies
5 Limitations and Future Works
Despite its competitive GAR accuracy and input flexibility, PromptGAR faces limitations with input resolution. (a) The input resolution is restricted to $224\times 224$, which is significantly lower than $1080\times 720$ in other methods [50, 29, 19], due to MViTv2’s 3D convolutions and correponding CUDA memory limitations. (b) In order to prevent out-of-memory OOM errors, only the smallest one from the MViTv2’s multi-scale feature maps ($8\times 7\times 7$ in the Volleyball dataset) can be utilized.
For future works, our research will explore several avenues to further enhance and validate PromptGAR’s capabilities. Firstly, given its support for single-actor prompts, we aim to assess its generalization capabilities on challenging action recognition tasks. Furthermore, to extend the experiments on multi-group GAR beyond the synthetic video mosaic dataset, we plan to evaluate PromptGAR on more complex datasets like MEVA [10], which features larger spatial movement, smaller actors, and human-vehicle group activities.
6 Conclusion
In this work, we present PromptGAR, a novel framework that addresses the limitations of current Group Activity Recognition (GAR) approaches by leveraging multi-modal prompts to achieve both input flexibility and high recognition accuracy. Extensive experiments show that PromptGAR allows the input flexibility across prompts, frames, and instances without the need for retraining. Furthermore, we introduce a novel relative instance attention mechanism that directly encodes instance IDs, ensuring long-term consistency for extended activity durations. Finally, PromptGAR also uses the area prompts to enable the selective recognition of the particular group activity among multiple concurrent groups. Extensive evaluation validates the effectiveness and generalizability of our model.
Supplementary Material
1 More on Implementation Details
Data Augmentations. Following MViTv2 in action recognition [31], we apply a comprehensive set of augmentations for group activity recognition (GAR): Random Augment [11], Random Resized Crop, Flip, Random Erasing [54], CutMix [51], and MixUp [53]. In the multi-group activity recognition (Multi-GAR) task, we remove the CutMix and MixUp, because it would be impossible to determine which specific group from the reference frame is being overlaid onto the target frame within the same area prompt.
$O$ Memory Top1 Mean
@@ 56 OOM - -
Method Modalities Head Inputs Re-train Top1 Mean RGB Bbox Kpt InstID $\widehat{X}_{gar}$ $\widehat{F}_{p}$ PromptGAR 96.0 96.3 95.7 95.8 94.1 94.3 95.8 96.0 95.5 95.7 94.0 94.2 93.9 94.4 95.2 95.6 95.2 95.6 93.4 94.0
@@ 2 Additional Ablation Studies
Head Inputs Tab. 10 indicates demonstrates the significance of prompt token inputs within the head, even when dealing with multi-modal prompts. We compared performance using a model trained on full prompts, varying the head inputs during testing. Under full modality inputs, using only the GAR class token $\widehat{\textbf{X}}_{gar}$ as the head input resulted in a minor 0.2% performance drop, while using only the prompt tokens $\widehat{\textbf{F}}_{p}$ led to a 0.8% drop. This shows that both inputs are crucial for PromptGAR’s high performance. Even with reduced modality inputs, both $\widehat{\textbf{X}}_{gar}$ only and $\widehat{\textbf{F}}_{p}$ only still yielded reasonable results. Notably, when using only prompt tokens $\widehat{\textbf{F}}_{p}$, the performance of full prompt inputs was similar to the one with skeleton-only inputs. This suggests that when RGB features don’t directly influence the final prediction, bounding box information is effectively captured within the skeleton data. In summary, this table showcases the flexibility of head inputs across various modalities and underscores the importance of both the RGB feature representation $\widehat{\textbf{X}}_{gar}$ and the prompt feature representation $\widehat{\textbf{F}}_{p}$ for optimal performance.
@@
Figure 8: The 2-Mosaic Volleyball Confusion Matrix
Figure 9: The 4-Mosaic Volleyball Confusion Matrix
@@ 3 Analysis
Quantitative Results. Fig. 7 - 9 shows the confusion matrix of Volleyball, NBA, 2-Mosaic Volleyball, and 4-Mosaic Volleyball respectively. (a) For Volleyball, while PromptGAR demonstrates high accuracy on the Volleyball dataset, a common misclassification occurs where ‘right-set’ is mistaken for ‘right-pass’. This error is attributed to the inherent similarity in player actions and positions, echoing findings in previous research [55]. (b) For NBA, errors highlight the challenge of distinguishing between highly similar actions. Specifically, differentiating ‘offensive’ and ‘defensive’ requires nuanced visual analysis of player uniform differences and rebound outcomes. Similarly, distinguishing ‘2p-fail’ from ‘2p-layup-fail’ demands precise attention to shooting position and posture near the rim. (c) In the 2-Mosaic and 4-Mosaic Volleyball datasets, PromptGAR achieves respectable results, despite the significant challenge of reduced actor size and motion blur. This difficulty arises because, while the input image size is doubled (or quadrupled), the model’s input resolution remains fixed at $224x\times 224$. Given that only area prompts are provided, without any additional enhancement information, the model struggles to distinguish between ‘pass’ and ‘set’ activities.

### v2

@@ 4.1 Experimental Setup
Prompt Types Top1 Acc Mean Acc Method RGB Bbox Kpt Flow Ball ARG [38] 90.7 91.0 HiGCIN [41] 91.4 92.0 DIN [45] $92.7^{\star}$ $92.8^{\star}$ POGARS [37] 93.9 - Composer [51] $93.6^{\star}$ - SkeleTR [11] 94.4 - MP-GCN [26] $^{\dagger} 95.095.0^{\dagger}$ Bi-Causal [49] $95.8^{\star}$ - CRM [3] 93.0 - ActorFormer [13] 94.4 - GIRN [31] 94.0 - SACRF [32] 95.0 - GroupFormer [24] 95.7 - Dual-AI [16] 95.4 - KRGFormer [30] 94.6 94.8 PromptGAR 96.0 96.3
Table 1: Quantitative Comparisons in Volleyball Dataset. (-) denotes not reported in the paper, (*) reproduced from released codes, ($\dagger$) reproduced with keypoint-ball results (late-fusion parts un-released), and unmarked other methods are not open-sourced.
Prompt Types Top1 Acc Mean Acc Method RGB Bbox Kpt Flow Ball ARG [38] 59.0 56.8 ActorFormer [13] 47.1 41.5 SAM [42] 49.1 47.5 DIN [45] 61.6 56.0 Dual-AI [16] 58.1 50.2 KRGFormer [30] 72.4 67.1 MP-GCN [26] $75.8^{\dagger}$ $72.0^{\dagger}$ DFWSGAR [20] 75.8 71.2 Flaming-Net [29] 79.1 76.0 PromptGAR 80.6 76.9
@@
Volleyball Dataset [18] contains 3,493 clips for training and 1,337 clips for testing. Each clip has 41 frames and is labeled with one of eight group activities. Bounding box and skeleton annotations, provided separately by [18, 51], are available only for the central 16 frames of each clip. Consistent with prior works, we limit our analysis for central 16 frames to avoid potential interference from additional group activities present in the remaining half of clips.
NBA Dataset [42] includes 7,624 training clips and 1,548 testing clips. Each clip includes 72 frames and is categorized into nine group activity labels. Due to its increasing complexity, the NBA dataset demands special design and processing compared to the Volleyball dataset [18]. Unlike Volleyball [18], we use all 72 frames in NBA because key event frames are not centrally located. Also, NBA activities are much longer than Volleyball’s. For example, ‘3p-fail-offensive’ requires recognizing shooter location, ball trajectory, and possession.
@@ 4.2 Group Activity Recognition
Volleyball Dataset. In Tab. 1, we compare our method to the group activity recognition methods that rely on: (a) RGB frames, like DIN [45], which suffers from limited performance; (b) skeletons and ball trajectories, like Composer [51] and MP-GCN [26], which requires ball trajectories as extra inputs; (c) combined visual prompts, like GroupFormer [24] and Dual-AI [16], which takes computationally expensive optical flows as inputs and thus sensitive to input frame rates. In contrast, our PromptGAR, without the drawbacks of those other visual prompt inputs, still produces 96.0% top1 accuracy and 96.3% mean accuracy, achieving considerable improvements across different baseline models. Also, ACCG [39] can not be evaluated due to unreleased codes.
NBA Dataset. Tab. 2 shows that PromptGAR gets competitive performance over previous methods on the challenging NBA dataset [42]. Prior approaches often face limitations. Older models like SAM [42] and Dual-AI [16] accept only raw videos and bounding boxes, which restricts their performance; Other methods, such as MP-GCN [26], require specialized inputs like ball trajectories, thereby limiting their applicability to broader, non-sports GAR tasks. Additionally, weakly-supervised techniques like DFWSGAR [20] and Flaming-Net [29] introduce considerable training complexity. In contrast, our PromptGAR, without these drawbacks, still produces 80.6% top1 accuracy and 76.9% mean accuracy, reflecting substantial advancements over various baseline models.
4.3 Input Flexibility
Prompt Types Method RGB Bbox Kpt Flow Re-train Top1 Acc Mean Acc 95.7 - 94.9 - GroupFormer [24] 94.1 - 96.0 96.3 95.7 95.8 94.1 94.3 PromptGAR 93.9 94.4
@@ 4.3 Input Flexibility
4.4 Actor Consistency
Method Shuffle Input Actor Order Re-train Top1 Acc Mean Acc MP-GCN [26] 75.8 72.0 74.7 71.2 PromptGAR 80.6 76.9 80.6 76.9
Table 6: Actor Order Invariance. For NBA dataset, input actor order is randomly shuffled, and results are averaged over five trials.
Method Relative Instance Identity Embeddings Re-train Top1 Acc Mean Acc PromptGAR 96.0 96.3 93.8 94.1
Table 7: Necessity of Instance Identities. In Volleyball dataset, relative instance identity embeddings are included or omitted, using prompts from bounding boxes, skeletons, and RGB videos.
Actor Order Invariance. Prompt inputs for the testing dataset are formatted as a tensor of shape $(N,T,M,J,C)$, representing the number of videos, frames, actors, point types, and coordinates, respectively. Bounding boxes and skeletal keypoints belonging to the same instance ID share the same index along the $M$ (actor) axis. The actor order along this $M$ axis is usually fixed in conventional evaluations, but we introduce a rigorous test: we randomly shuffle the $M$ axis and evaluate the same model. This procedure is repeated five times, and the averaged performance is presented in Tab. 6. Our results demonstrate that after shuffling, MP-GCN [26] shows a slight performance degradation, but our model maintains identical performance numbers. This is because current GAR methods [51, 26, 13, 24, 16] fix that order for both validation and testing, choosing the best checkpoint based on performance with that specific player order. In contrast, our model incorporates a novel relative instance attention mechanism that encodes only whether two actors are the same or not. This design ensures our model remains unaffected by actor order in real-world scenarios, thereby enhancing its robustness.
Necessity of Instance Identities. As shown in Tab. 7, there is a performance drop without instance identities. This is because, without instance identities, the model struggles to maintain actor consistency and reliably associate individual features across different frames. This also proves the necessity of our instance identities.
@@ 4.4 Actor Consistency
$O$ Memory Top1 Acc Mean Acc
@@ 56 OOM - -
4.5 Ablation Studies
Tab. 8 details the impact of our novel modules on performance, starting with a baseline of 96.0% top-1 accuracy on the Volleyball dataset. (a) Relative instance attention. Eliminating instance IDs during training and testing resulted in a 0.6% drop in top-1 accuracy. This confirms that instance IDs are vital for encoding long-term consistency, enhancing performance. Furthermore, relative instance attention also accelerated training convergence. (b) Prompt tokens in head. Removing prompt tokens $\widehat{\textbf{F}}_{p}$ from the head, leaving only the GAR class token $\widehat{\textbf{X}}_{gar}$, led to a 0.3% accuracy decrease. This highlights the necessity of explicitly incorporating prompt tokens. Without them, the model struggles to integrate diverse visual prompt features, as the GAR class token alone sufficiently captures RGB information. (c) No prompts. Disabling the entire prompt encoder, using only the GAR class token $\widehat{\textbf{X}}_{gar}$ as input for the recognition decoder and the head, further reduced accuracy by 0.8%. This underscores the contribution of visual prompt inputs in PromptGAR. Crucially, even without any prompts during training and testing, our model still compares favorably with other prompt-based baseline methods. Specifically, it achieves a 1.6% higher accuracy than DIN, which utilizes both RGB frames and bounding boxes, and a 0.7% higher accuracy than Composer, which uses skeletons and ball trajectories as additional information. This demonstrates the superior ability of our MViTv2 architecture to process spatial-temporal information within only the RGB features compared to other backbones that process the video frame-by-frame separately.
5 Limitations
Despite its competitive GAR accuracy and input flexibility, PromptGAR faces limitations with input resolution. The input resolution is restricted to $224\times 224$, which is significantly lower than $1080\times 720$ in other methods [45, 24, 16].
6 Conclusion
In this work, we present PromptGAR, a novel framework that addresses the limitations of current Group Activity Recognition (GAR) approaches by leveraging diverse visual prompts to achieve both input flexibility and high recognition accuracy. Extensive experiments show that PromptGAR allows the input flexibility across prompts, frames, and instances without the need for retraining. Furthermore, we introduce a novel relative instance attention mechanism that directly encodes instance identities, ensuring actor consistency. Extensive evaluation validates the effectiveness and applicability of our model.
Supplementary Material
Method Prompt Types Head Inputs Re-train Top1 Mean RGB Bbox Kpt InstID $\widehat{X}_{gar}$ $\widehat{F}_{p}$ PromptGAR 96.0 96.3 95.7 95.8 94.1 94.3 95.8 96.0 95.5 95.7 94.0 94.2 93.9 94.4 95.2 95.6 95.2 95.6 93.4 94.0
@@ 6 Conclusion
Train Test Method RGB Bbox Kpt InstID RGB Bbox Kpt InstID Top1 Acc Mean Acc 96.0 96.3 95.7 95.8 95.4 95.8 PromptGAR 94.3 94.7
Table 12: Low Reliance on Skeletons.
1 More on Implementation Details
Data Augmentations. Following MViTv2 in action recognition [25], we apply a comprehensive set of augmentations for group activity recognition (GAR): Random Augment [9], Random Resized Crop, Flip, Random Erasing [50], CutMix [46], and MixUp [48].
@@ 2 Additional Ablation Studies
Head Inputs. Tab. 11 indicates demonstrates the significance of prompt token inputs within the head, even when dealing with diverse visual prompts. We compared performance using a model trained on full prompts, varying the head inputs during testing. Under full prompt inputs, using only the GAR class token $\widehat{\textbf{X}}_{gar}$ as the head input resulted in a minor 0.2% performance drop, while using only the prompt tokens $\widehat{\textbf{F}}_{p}$ led to a 0.8% drop. This shows that both inputs are crucial for PromptGAR’s high performance. Even with reduced prompt inputs, both $\widehat{\textbf{X}}_{gar}$ only and $\widehat{\textbf{F}}_{p}$ only still yielded reasonable results. Notably, when using only prompt tokens $\widehat{\textbf{F}}_{p}$, the performance of full prompt inputs was similar to the one with skeleton-only inputs. This suggests that when RGB features don’t directly influence the final prediction, bounding box information is effectively captured within the skeleton data. In summary, this table showcases the flexibility of head inputs across various visual prompts and underscores the importance of both the RGB feature representation $\widehat{\textbf{X}}_{gar}$ and the prompt feature representation $\widehat{\textbf{F}}_{p}$ for optimal performance.
Low Reliance on Skeletons. Tab. 12, our model has competitive performance even when trained without skeletal data. This outcome indicates that our model does not heavily rely on skeletal information.
@@ 3 Analysis
Quantitative Results. Fig. 7 - 7 shows the confusion matrix of Volleyball and NBA, respectively. (a) For Volleyball, while PromptGAR demonstrates high accuracy on the Volleyball dataset, a common misclassification occurs where ‘right-set’ is mistaken for ‘right-pass’. This error is attributed to the inherent similarity in player actions and positions, echoing findings in previous research [51]. (b) For NBA, errors highlight the challenge of distinguishing between highly similar actions. Specifically, differentiating ‘offensive’ and ‘defensive’ requires nuanced visual analysis of player uniform differences and rebound outcomes. Similarly, distinguishing ‘2p-fail’ from ‘2p-layup-fail’ demands attention to shooting position and posture near the rim.